# On the Need to Rethink Trust in AI Assistants for Software Development: A Critical Review

### v2

@@ IV Adopting Trust Models in SE (RQ3)
To answer RQ3, we first discuss problems with the way SE research currently handles trust research and then describe how the common trust model of [95] [95] could be adopted to study trust in AI assistants for software development.
@@ IV-A Current State of Trust Research in SE
The detachment of SE research from established trust models is problematic, as evidenced by the equation of trust with the acceptance of generated content, which is shortsighted for several reasons. Drawing on the foundations of trust from psychology and philosophy that we introduced in Section II, Table III shows how such an operationalization disregards the tenets of trust discussed in these disciplines. Most importantly, both disciplines see trust as an attitude and not a behavior. Therefore, measuring acceptance disregards factors that influence a person’s behavior, such as their situation (e.g., time pressure), the type of AI system they are interacting with, their specific task, the potential consequences of decisions, and many more. Still, even if acceptance measured trust, it would not measure calibrated or warranted trust. Hence, the measured trust could be completely detached from the actual capabilities of an AI assistant. However, given the fact that GenAI-based software development assistants will continue to hallucinate [14], the potentially resulting maintainability challenges [87], security risks [111], and negative impact on learning [162, 131], it would not only be morally but also pragmatically better to find out whether developers’ trust in such tools is actually calibrated or warranted. After all, maintainability issues or security vulnerabilities introduced by relying on generated code could cause major financial and reputational damage to individuals and companies.
@@
It was interesting for us to see that IS and HCI primarily focus on the psychological side of trust. We posit that discussion in SE can also benefit from incorporating philosophical insights on trust. For example, the discussion as to whether trust in AI is even conceptually possible is relevant in the sense that it critically questions whether engineering for trust is not just a marketing phenomenon or, in the worst case, even ethicswashing (see Section II-C). A similar phenomenon can be observed in the debates on XAI, where the expectation is that XAI invariably leads to greater trust, something that could not be proven in studies [70]. In this sense, the concept of justification is also important for SE, as it can help ensure that trust in AI assistants for software development is formed appropriately, rather than, for example, being the result of misleading marketing rhetoric. It is surprising that the trust conceptualizations in SE are still so underdeveloped, as tools from SE can be extremely important for creating trustworthy systems. For example, [2] argue that SE processes and tools, in particular a detailed requirements analysis, monitoring instruments, and automated testing, are essential to create systems that can really be trusted [2].
IV-B An Example of Adopting Trust Models in SE
@@
In general, the dimensions of perceived trustworthiness proposed by [95] [95], that is, ability (A), benevolence (B), and integrity (I) can be easily transferred to AI assistants for software development. As mentioned before, ability requires AI assistants to generate correct and reliable code that is appropriate for a given context. Benevolence reflects interaction patterns that AI vendors implement to ensure that AI assistants are perceived as supportive, and integrity models the data curation and guardrails that AI vendors and regulatory bodies implement to avoid biases or problematic code that, for example, contains security vulnerabilities. In the following, we discuss existing instruments that SE researchers can use, with a particular focus on our own adaption of [95]’s trust model as presented in Figure 1. Please note that this is one example for adapting an established trust model to an SE context. If concepts such as calibrated trust are important for a particular SE research project, the model by [80] might be more suitable [80]. It is important that SE researchers know enough about trust conceptualizations to understand the implications of adapting particular trust models themselves. We hope that our article serves as a starting point for this.
@@
As already mentioned, operationalizing trust is difficult; one needs to be specific about the actual concepts that one intends to capture (e.g., the perceived trustworthiness of a tool or a person’s general propensity to trust). Fortunately, validated instruments exist to assess [95]’s ABI dimensions of perceived trustworthiness. SE researchers can adapt these instruments to study trust in AI assistants [94, 99]. There further exist various instruments to assess a person’s propensity to trust [161, 99, 43, 128]. Since perceived risk and vulnerability are central to [95]’s model, we also want to refer to instruments designed to measure related concepts such as a person’s willingness to be vulnerable or their trusting intensions [129, 99, 48].
@@
As mentioned above, IS literature has discussed operationalizing trust based on observed behavior by calculating the probability that users follow a model’s prediction based on usage data [115, 125]. Since this is closely related to our motivating example of software developers accepting code suggestions, we want to discuss how such metrics can be included in our representation of [95]’s trust model (Figure 1). [126] argue that behavioral metrics might work well as as proxies of trust when comparing different treatments (e.g., different ways of presenting AI suggestions) or comparing different groups based on the same treatment [126]. However, factors such as participant knowledge and confidence, and (perceived) task difficulty are known to affect advice acceptance [164]. Over time, the usage of tools moderates users’ bias towards self-reliance in decision-making and could further influence the observed behavior [39].
@@
In the context of [95]’s model, the acceptance rate for AI suggestions can be interpreted as an indicator of the tool’s “observed outcome.” Studying such data in detail is crucial, as acceptance of AI suggestions can be a factor by which people form their trust; that is, in [95]’s model, acceptance rates can contribute to the perceived ability of the system. Support for this view comes, for example, from [154]’s study. They have developed and tested design concepts, including a dashboard showing a user’s personal GitHub Copilot acceptance rate, finding that it helped developers align their expectations with an AI assistant’s ability [154].
@@
A general limitation of the questionnaire-based instruments referenced above is that answering survey questions can never fully capture trust as conceptualized by [95]. They cannot capture critical situations in which a person actually made itself vulnerable to the actions of another party [95] (here: an AI assistant). As mentioned above, proxies of behavioral trust can complement questionnaire-based instruments. However, also they also cannot capture actual behavioral trust. An general example of a situation in which behavioral trust occurs would be that parents leave their children with a certain babysitter as an indicator of trusting the babysitter [126]. Another limitation is that human behavior and decision-making are complex and can be affected by factors such as monetary constraints or a lack of alternatives. Therefore, researchers must be careful when using behavior as a proxy for trust, considering that there might be different drivers for the observed behavior.
IV-C Recommendations
@@
    Before designing concrete studies on trust an AI assistants, carefully read the literature on trust beyond SE. We hope that our article serves as a useful starting point for this.
@@ V Threats to Validity
As with any empirical study, our methodology is subject to certain limitations that may affect the completeness, accuracy, and generalizability of our findings. In the following, we identify and discuss potential threats to the validity of our methodology, including biases in paper selection, interpretation of trust conceptualizations, and the applicability of findings across disciplines. We classify the validity threats in three categories as proposed by [159] [159].
Internal Validity concerns whether our study design and execution introduced biases that may affect the correctness of our findings. A potential threat to internal validity is that we may have missed indirect or implicit definitions of trust as well as references to trust models during our literature review. Additionally, by focusing on the highest-ranked articles from our queries, we may have unintentionally excluded relevant papers ranked lower. We acknowledge the selective inclusion based on an opaque ranking algorithm provided by Google Scholar as a limitation. However, Google Scholar is used by academics across disciplines and allowed us to use the same sampling approach for all disciplines. Future work should extend our literature review by using discipline-specific retrieval approaches. In addition, we complemented the review of SE literature by conducting a critical review of recently published SE papers, which confirmed the results of our review based on Google Scholar search results.

### v3

@@ IV Adopting Trust Models in SE (RQ3)
To answer RQ3, we first discuss problems with the way SE research currently handles trust research and then describe how the common trust model of [Mayer1995] [Mayer1995] could be adopted to study trust in AI assistants for software development.
@@ IV-A Current State of Trust Research in SE
The detachment of SE research from established trust models is problematic, as evidenced by the equation of trust with the acceptance of generated content, which is shortsighted for several reasons. Drawing on the foundations of trust from psychology and philosophy that we introduced in Section II, Table III shows how such an operationalization disregards the tenets of trust discussed in these disciplines. Most importantly, both disciplines see trust as an attitude and not a behavior. Therefore, measuring acceptance disregards factors that influence a person’s behavior, such as their situation (e.g., time pressure), the type of AI system they are interacting with, their specific task, the potential consequences of decisions, and many more. Still, even if acceptance measured trust, it would not measure calibrated or warranted trust. Hence, the measured trust could be completely detached from the actual capabilities of an AI assistant. However, given the fact that GenAI-based software development assistants will continue to hallucinate [DBLP:journals/corr/abs-2409-05746], the potentially resulting maintainability challenges [DBLP:journals/tosem/LiuLWTLLL24], security risks [DBLP:conf/ccs/PerryS0B23, DBLP:conf/sp/PearceA0DK22, DBLP:conf/uss/SandovalPNKGD23, DBLP:conf/chi/SerafiniYN25, DBLP:conf/sac/KudriavtsevaHG25, DBLP:conf/wcre/MajdinasabBRDTK24, DBLP:conf/ccs/KlemmerHPLBPMRV24, DBLP:conf/sp/OhLPKK24], and negative impact on learning [14, DBLP:journals/cacm/Shein24b], it would not only be morally but also pragmatically better to find out whether developers’ trust in such tools is actually calibrated or warranted. After all, maintainability issues or security vulnerabilities introduced by relying on generated code could cause major financial and reputational damage to individuals and companies.
@@
It was interesting for us to see that IS and HCI primarily focus on the psychological side of trust. We posit that discussion in SE can also benefit from incorporating philosophical insights on trust. For example, the discussion whether trust in AI is even conceptually possible is relevant in the sense that it critically questions whether engineering for trust is not just a marketing phenomenon or, in the worst case, even ethicswashing (see Section II-C). A similar phenomenon can be observed in the debates on XAI, where the expectation is that XAI invariably leads to greater trust, something that could not be proven in studies [Kaestner2021Relation]. In this sense, the concept of justification is also important for SE, as it can help ensure that trust in AI assistants for software development is formed appropriately, rather than, for example, being the result of misleading marketing rhetoric. It is surprising that the trust conceptualizations in SE are still so underdeveloped, as tools from SE can be extremely important for creating trustworthy systems. For example, [DBLP:conf/ecai/AhujaBBCGLMSSS20] argue that SE processes and tools, in particular a detailed requirements analysis, monitoring instruments, and automated testing, are essential to create systems that can really be trusted [DBLP:conf/ecai/AhujaBBCGLMSSS20].
IV-B General Role of Trust in SE
In software development, negative outcomes, such as bugs or security vulnerabilities introduced by using AI assistants, materialize with a certain delay, and are often borne by someone else. Human reviewers must invest more time reviewing large amounts of generated code, security teams must catch vulnerabilities in production, bugs slow users down, and the organization may face a reputation loss due to data leakage or outages. This creates an incentive landscape in which an individual developer can rationally feel a low personal risk even when the system-level risk is unchanged or even increased. In trust terms, this can lead to an unjustifiably low perceived risk and potentially even to an unjustifiably high trust propensity.
Since developers are rewarded for throughput (e.g., resolved tickets), and costs of defects accumulate as technical debt, teams might drift into thinking (and trusting) that “it usually works.”—while long-term maintenance costs are accumulating. This makes trust in software contexts unusually sensitive to incentives, governance, and feedback frequency. Such factors are different in settings where the consequences are immediate and personal. In SE, however, a developer might have already left the company before the problems, which their trust in AI assistants has caused, materialize.
Moreover, the object that is trusted can shift in an SE context. Software development is embedded in institutionalized safety nets that involve code reviews, software tests, and compliance checks, often automated or supported by continuous integration pipelines. These processes can be risk-reducing, so developers may come to trust the pipeline rather than the AI assistant (our tests will catch potential issues). This mindset can lead to trust laundering, where weak trust in an AI output is converted into or substituted by greater trust in the infrastructure. The effect of this is that trust becomes procedural and distributed. It is not I trust the AI assistant but I trust our socio-technical infrastructure to detect and absorb errors.
Another philosophical and legal issue is that AI assistants lead to ambiguous authorship and accountability. When AI suggestions are integrated, it can become even more unclear who is responsible for an issue: Is it the developer who used the AI assistant, the reviewer who approved the change, or the team whose tests did not catch the bug? If accountability becomes more diffuse, developers may feel less personally exposed, which again raises trust even though organizational risk remains the same. In contrast, in teams with strong code ownership, norms, and guidelines, the same ambiguity can reduce trust because accepting AI output threatens clear boundaries of responsibility. One could argue that trust really depends on the institutional practices in place.
IV-C An Example of Adopting Trust Models in SE
@@
The dimensions of perceived trustworthiness proposed by [Mayer1995] [Mayer1995], that is, ability (A), benevolence (B), and integrity (I) can be easily transferred to AI assistants for software development. As mentioned before, ability requires AI assistants to generate correct and reliable code that is appropriate for a given context. Benevolence reflects interaction patterns that AI vendors implement to ensure that AI assistants are perceived as supportive, and integrity models the data curation and guardrails that AI vendors and regulatory bodies implement to avoid biases or problematic code that, for example, contains security vulnerabilities. In the following, we discuss existing instruments that SE researchers can use, with a particular focus on our own adaption of [Mayer1995]’s trust model as presented in Figure 1. Please note that this is one example for adapting an established trust model to an SE context. If concepts such as calibrated trust are important for a particular SE research project, the model by [Lee2004Trust] might be more suitable [Lee2004Trust]. It is important that SE researchers know enough about trust conceptualizations to understand the implications of adapting particular trust models themselves. We hope that our article serves as a starting point for this.
@@
As already mentioned, operationalizing trust is difficult; one needs to be specific about the actual concepts that one intends to capture (e.g., the perceived trustworthiness of a tool or a person’s general propensity to trust). Fortunately, validated instruments exist to assess [Mayer1995]’s ABI dimensions of perceived trustworthiness. SE researchers can adapt these instruments to study trust in AI assistants [MayerDavies1999, DBLP:journals/isr/McKnightCK02]. There further exist various instruments to assess a person’s propensity to trust [YamagishiYamagishi1994, DBLP:journals/isr/McKnightCK02, Frazier01102013, Scholz17012025]. Since perceived risk and vulnerability are central to [Mayer1995]’s model, we also want to refer to instruments designed to measure related concepts such as a person’s willingness to be vulnerable or their trusting intensions [DavidSchoorman02012016, DBLP:journals/isr/McKnightCK02, gillespie201520].
@@
As mentioned above, IS literature has discussed operationalizing trust based on observed behavior by calculating the probability that users follow a model’s prediction based on usage data [DBLP:conf/chi/Poursabzi-Sangdeh21, DBLP:journals/corr/abs-1901-08558]. Since this is closely related to our motivating example of software developers accepting code suggestions, we want to discuss how such metrics can be included in our representation of [Mayer1995]’s trust model (Figure 1). [DBLP:journals/jds/SchmidtBT20] argue that behavioral metrics might work well as proxies of trust when comparing different treatments (e.g., different ways of presenting AI suggestions) or comparing different groups based on the same treatment [DBLP:journals/jds/SchmidtBT20]. However, factors such as participant knowledge and confidence, and (perceived) task difficulty are known to affect advice acceptance [YANIV20041]. Over time, the usage of tools moderates users’ bias toward self-reliance in decision-making and could further influence the observed behavior [DBLP:journals/ijmms/DzindoletPPPB03].
@@
In the context of [Mayer1995]’s model, the acceptance rate for AI suggestions can be interpreted as an indicator of the tool’s “observed outcome.” Studying such data in detail is crucial, as acceptance of AI suggestions can be a factor by which people form their trust; that is, in [Mayer1995]’s model, acceptance rates can contribute to the perceived ability of the system. Support for this view comes, for example, from [12]’s study. They have developed and tested design concepts, including a dashboard showing a user’s personal GitHub Copilot acceptance rate, finding that it helped developers align their expectations with an AI assistant’s ability [12].
@@
A general limitation of the questionnaire-based instruments referenced above is that answering survey questions can never fully capture trust as conceptualized by [Mayer1995]. They cannot capture critical situations in which a person actually made itself vulnerable to the actions of another party [Mayer1995] (here: an AI assistant). As mentioned above, proxies of behavioral trust can complement questionnaire-based instruments. However, they also cannot capture actual behavioral trust. A general example of a situation in which behavioral trust occurs would be that parents leave their children with a certain babysitter as an indicator of trusting the babysitter [DBLP:journals/jds/SchmidtBT20]. Another limitation is that human behavior and decision-making are complex and can be affected by factors such as monetary constraints or a lack of alternatives. Therefore, researchers must be careful when using behavior as a proxy for trust, considering that there might be different drivers for the observed behavior.
IV-D Recommendations
@@
    Before designing concrete studies on trust in AI assistants, carefully read the literature on trust beyond SE. We hope that our article serves as a useful starting point for this.
@@
{results}
RQ3 (Recommendations for SE):
SE researchers should treat trust as a multidimensional and context-dependent phenomenon. Focusing on individual factors is generally not sufficient to holistically capture trust. Therefore, SE researchers should build on established trust models and measurement instruments. Section IV-D provides further recommendations.
@@ V Threats to Validity
As with any empirical study, our methodology is subject to certain limitations that may affect the completeness, accuracy, and generalizability of our findings. In the following, we identify and discuss potential threats to the validity of our methodology, including biases in paper selection, interpretation of trust conceptualizations, and the applicability of findings across disciplines. We classify the validity threats in three categories as proposed by [wohlin2012experimentation] [wohlin2012experimentation].
Internal Validity concerns whether our study design and execution introduced biases that may affect the correctness of our findings. A potential threat to internal validity is that we may have missed indirect or implicit definitions of trust as well as references to trust models during our literature review. Additionally, by focusing on the highest-ranked articles from our queries, we may have unintentionally excluded relevant papers ranked lower. We acknowledge the selective inclusion based on an opaque ranking algorithm provided by Google Scholar as a limitation. However, Google Scholar is used by academics across disciplines and allowed us to use the same sampling approach for all disciplines. We complemented the review of SE literature based on Google Scholar search results by conducting a critical review of recently published SE papers, confirming our earlier results.
Future work should extend the literature review by using discipline-specific retrieval approaches. One related discipline that studies AI assistants for software development [DBLP:conf/ccs/PerryS0B23, DBLP:conf/sp/PearceA0DK22, DBLP:conf/uss/SandovalPNKGD23], sometimes referring to human trust [DBLP:conf/sp/OhLPKK24, DBLP:conf/ccs/KlemmerHPLBPMRV24], is the security research community. We did not explicitly exclude security-related research in our review, only mentions of trust in a more technical context, as clarified in Section I. However, at the same time, we did not explicitly include security research venues. Extending the literature review to consider security research as well is a promising direction for future work.